OpenStudy (anonymous):

A person standing close to a railroad crossing hears the whistle of an approaching train. He notes that the pitch of the whistle drops as the train passes by and moves away from the crossing. The frequency of the approaching whistle is 524 Hz, and drops to 497 Hz after the train is well past the crossing. What is the speed of the train? Use 340 m/s for the speed of sound in air.

OpenStudy (anonymous):

its the dobbler effect

OpenStudy (anonymous):

Yes it is.

OpenStudy (anonymous):

We know that the frequency of a wave is related to wavelength and velocity as such. \[f = {v \over \lambda}\]

OpenStudy (anonymous):

"Dobbler" indeed. I'm sorry, I just found that amusing.

OpenStudy (anonymous):

its f=(v-V/v-V)*fo

OpenStudy (anonymous):

v=340m/s but i am not sure what to do after that

OpenStudy (anonymous):

As the train moves towards us, the relative velocity of the wave is increased. As it moves away, the relative velocity decreases. This creates a frequency shift.

OpenStudy (anonymous):

so when it comes to us its v+V and when it goes away its v-V?

OpenStudy (anonymous):

Slow down. Let's work through it.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The doppler effect is expressed as\[f = \left ( c + v_r \over c + v_s \right)f_o\]In this case, we are standing still, therefore \(v_r = 0\). The doppler effect equation simplifies to \[f = \left ( c \over c + v_s \right) f_o\]where \(v_s\) is the velocity of the train.

OpenStudy (anonymous):

We know the frequency as the train approaches and as it leaves, and we know the velocity of the sound wave. We also know that \(f_o\) is constant. We can write two expressions of the doppler effect equation. One as the train approaches and one as the train leaves. Moving toward observer:\[f_1 = \left ( c \over c - v_s \right) f_o\]Moving away from observer:\[f_2 = \left ( c \over c + v_s \right ) f_o\]

OpenStudy (anonymous):

Let's rearrange both equations to be in terms of \(f_o\). \[f_o = f_1 \left ( c - v_s \over c \right )\]and\[f_o = f_2 \left ( c + v_s \over c \right)\] Now, let's set them equal to each other.

OpenStudy (anonymous):

ok, is fo the initial? so 544?

OpenStudy (anonymous):

*524

OpenStudy (anonymous):

\[f_1 \left( c - v_s \over c \right) = f_2 \left( c+ v_s \over c \right)\]

OpenStudy (anonymous):

\(f_1 = 524 Hz\) and \(f_2 = 497 Hz\).

OpenStudy (anonymous):

We can solve that equation for \(v_s\)

OpenStudy (anonymous):

can u explain to me what vs and vr mean?

OpenStudy (anonymous):

\(v_s\) is the speed of the source. In this case, it is the speed of the train. \(v_r\) is the speed of the observer. In this case, \(v_r = 0\)

OpenStudy (anonymous):

oh ok, i found vs to be 340 m/s :S?

OpenStudy (anonymous):

oh wait hold on

OpenStudy (anonymous):

Nope. If there were the case, f would be infinity as the source approached and we wouldn't hear anything.

OpenStudy (anonymous):

ya cuz 340-340 i found an answer of 8.99m/s

OpenStudy (anonymous):

\[f_1 \left ( 1 - {v_s \over c} \right) = f_2 \left( 1 + {v_s \over c} \right)\]\[f_1 - f_2 = f_1 {v_s \over c} + f_2 {v_s \over c}\]\[f_1 - f_2= {v_s \over c} (f_1 + f_2)\]\[v_s = \left (f_1 - f_2 \over f_1 + f_2 \right) c\]

OpenStudy (anonymous):

yes i got it :D i think i kind of get the concept thank you so much :)

OpenStudy (anonymous):

I also got 8.99 m/s.